OpenStudy (anonymous):

Three plots having an area of 132, 204 and 228 square mts. Respectively are to be sub-divided into equalized flower beds. If the breadth of a bed is 3mts., find the maximum length that a bed.

OpenStudy (kc_kennylau):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (kc_kennylau):

That's the same as finding the GCD of \(\dfrac{132}3\), \(\dfrac{204}3\) and \(\dfrac{228}3\).

OpenStudy (anonymous):

33,68,76